ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

15. Numerical Analysis and Scientific Computing

1991 MS Classification: 65 Zhou, Aihui, Institute of Systems Science, Academia Sinica, China Multiparameter error resolution for finite elements It is proven that the error behaviour of finite element solutions could be resolved by multi-scale of the mesh. With introducing some suitable multiparameters of mesh sizes, a parallel algorithm for higher accuracy approximations is proposed, and a local refinement strategy is presented, too. This approach is mainly based on an error asymptotic resolution. Part of these results is joint work with my colleagues. Reference: A. Zhou, etal, A multi-parameter splitting extrapolation and a parallel algorithm, Sys. Sci. & Math. Sci. , 10, 3(1997), 253-260.